Teacher is throwing a party for students using $136. She plans on buying 5 Pizzas and 40 Drinks.
Previously it cost her 22.80 for 1 Pizza and 6 Drinks.
What do the 2 variables represent.
Write a system of equations to represent the model
What is the cost of 1 Drink.? What is the cost of 1 Pizza?

Answer by Hawksfan(61)

Let P = # of Pizzas
Let D = # of Drinks
We need to write two equations with two unknowns and solve the problem.
First equation:
5P + 40D = 136
Second equation :
1P + 6D = 22.80
Now combine the equations by putting them side by side:
5P + 40D = 136
1P + 6D = 22.80 we need to get the additive opposite of P or D
lets multiply the second equation by -5
5P + 40D =136
-5P+ -30D = -114 now add the two equations
10D = 22
D = 2.2 or 2 drinks
1P + 6D = 22.80
P + 6(2.2) = 22.80
P +13.2 = 22.8
P = 9.6 pizzas or 9 pizzas
you can also solve this by substitution